Find the sum of vector A and B
A=2i +2j
B=2i - 4j
1
Expert's answer
2020-02-24T10:10:22-0500
"\bold{A+B}=(2\bold{i}+2\bold{j})+(2\bold{i}-4\bold{j})"

"\bold{A+B}=(2+2)\bold{i}+(2-4)\bold{j}"

"\bold{A+B}=4\bold{i}-2\bold{j}"